OpenStudy (anonymous):

HELP When solving the equation ½t = -7, the next logical step is multiply both sides of the equation by 2. The justification is the

OpenStudy (thesecret20111):

Because multiplying 1/2 by 2 will give 1 That will make the eqaution t=-7*2 t=-14

OpenStudy (anonymous):

it's asking for a reasoning :/ Reflexive Property of Equality Symmetric Property of Equality Subtraction Property of Equality Multiplication Property of Equality

jimthompson5910 (jim_thompson5910):

It's a property of equality (since you're doing something to both sides of an equation)

jimthompson5910 (jim_thompson5910):

the question is what are you doing to both sides?

jimthompson5910 (jim_thompson5910):

it's not a trick question (well I hope it doesn't seem that way)

OpenStudy (anonymous):

reflexive?

OpenStudy (thesecret20111):

I think it is the Multiplication Property of Equality

jimthompson5910 (jim_thompson5910):

you're reflexing each side?

OpenStudy (anonymous):

no your multiplying :D

OpenStudy (anonymous):

by 2

jimthompson5910 (jim_thompson5910):

so it's the multiplicative property of equality

jimthompson5910 (jim_thompson5910):

that says if you have a = b then you can multiply both sides by c to get c*a = c*b

OpenStudy (anonymous):

thank you so much :D best helper i seen on this site :D

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

can you help me on one more it's a drawing :/

jimthompson5910 (jim_thompson5910):

alright go for it

OpenStudy (anonymous):

i think its definition :/ not 100 percent sure though

jimthompson5910 (jim_thompson5910):

nope it's not that and it's not the symmetric property either

jimthompson5910 (jim_thompson5910):

when you say "the measure of angle 1 + the measure of angle 2", you are basically adding the two angles because they are adjacent angles, they form a larger angle (think of them glued together)

jimthompson5910 (jim_thompson5910):

so which property am I describing?

OpenStudy (anonymous):

angle addition -___-

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

sorry for these really simple questions

jimthompson5910 (jim_thompson5910):

a lot of these have descriptive names so they are probably deceptively simple

jimthompson5910 (jim_thompson5910):

its fine, you're just second guessing yourself

OpenStudy (anonymous):

i think it might be segment addition

jimthompson5910 (jim_thompson5910):

no you're still adding angles

jimthompson5910 (jim_thompson5910):

not segments

OpenStudy (anonymous):

lol -___- transitive prop of congruence

jimthompson5910 (jim_thompson5910):

technically it should be "the definition of complementary angles", but the next best thing is "angle addition postulate" because you're basically exploiting the fact that the two add to 90 degrees

OpenStudy (e.mccormick):

@arkoffusion If you don't know and ask, it is fine. That is how you learn. It is if you don't ask that it is a problem! So don't apologize for simple questions. It is still an important part of learning!